3. In a city of the California, 12318 flu cases were identified in 2015. The total pomulation for the comty is 171.029 in the sidle year of 2015. The total number of death in this county in 2015 is 3478. Among them, 12 died from influenza. Please calculate the al1-cause mortality, influenza specific mortality and case fatality for influenza. (3 points).
Answer #1

All-cause mortality = (Number of deaths / Total population in middle of year) * 1000 = (3478 / 171029) * 1000 = 0.02033573 * 1000 = 20.34 per 1000

Influenza specific mortality = (Number of deaths due to influenza / Total population in middle of year) * 1000 = (12 / 171029) * 1000 = 0.07 per 1000

Case fatality for influenza = (Number of deaths due to influenza / Number of individuals with influenza disease) * 100

= (12 / 12318) * 100 = 0.097 %